The average hospitality girls working in a certain area is 19 years old. A civic oriented group made a recent survey on the age of the hospitality girls working in the same locality. A random sample of 20 respondents showed an average of 15.4 years old with a standard deviation of 2.14 years old. Using a 0.05 level of significance, has the hospitality girls gone lower?

1
Expert's answer
2021-12-21T18:30:11-0500

"H_0:\mu=19" , average hospitality girls working in a certain area is 19 years old

"H_a:\mu<19" , average hospitality girls working in a certain area is less than 19 years old


"t=\frac{\overline{x}-\mu}{\sigma/\sqrt n}=\frac{15.4-19}{2.14/\sqrt{20}}=-7.523"


"df=n-1=19"

critical value:

"t_{crit}=1.729"


Since "t<-1.729" we reject null hypothesis. Average hospitality girls working in a certain area is less than 19 years old.